OpenStudy (anonymous):

Find the derivative: h(y)=ln(y^2cosy) Please show steps, I'm completely lost.

sam (.sam.):

What's your progress?

OpenStudy (anonymous):

I have ln(y^2)+ln(cosy)

sam (.sam.):

We can do that by using product rule and some others

OpenStudy (anonymous):

OKay

sam (.sam.):

alright after the derivative do you get \[\large h'(y)=\frac{1}{x^2\cos(x)}(x^2(-\sin(x)+\cos(x)(2x))\]

sam (.sam.):

\[\large h'(y)=\frac{1}{x^2\cos(x)}(x^2(-\sin(x))+\cos(x)(2x))\]

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

\[y=\ln(u), \ where \ u \ is \ a \ function \ of \ x \ ,then\ y'=\frac{ 1 }{ u }\frac{ du}{ dx }\]

sam (.sam.):

You differentiate ln(x) you will get \(\huge \frac{1}{x}\), same for that, and change those "x" into y's I made a mistake instead of putting it "y"

sam (.sam.):

\[\large h'(y)=\frac{1}{y^2\cos(y)}(y^2(-\sin(y))+\cos(y)(2y))\]

OpenStudy (anonymous):

So the way I have it started is incorrect?

sam (.sam.):

Then after differentiating ln(y), you need to differentiate inside of it according to chain rule, you can do that by product rule

sam (.sam.):

You can do that, it doesn't matter, do you get \[h(y)=\ln(y^2)+\ln(\cos(y)) \\ \\ h'(y)=\frac{1}{y^2}(2y)+\frac{1}{\cos(y)}(-\sin(y))\]

OpenStudy (anonymous):

if f(x) =ln(x) then y' of f(x)=\[\frac{ 1 }{ x } \]

OpenStudy (anonymous):

Which way is the correct way?

OpenStudy (anonymous):

\[h'(y)=\frac{ 1 }{ y^2\cos(y)}\frac{ d }{ dy }y^2\cos(y)=\frac{ 1 }{ y^2 \cos(y)}2y \cos(y)-y^2\sin(y)=\frac{ 2\cos(y)-y \sin(y) }{ y \cos(y) }\]\[=\frac{ 2 }{ y }-\tan(y)\] @kaldrid3

OpenStudy (anonymous):

Do you see how that works?

OpenStudy (anonymous):

sam was right.

OpenStudy (anonymous):

I feel like h′(y)=1x2cos(x)(x2(−sin(x))+cos(x)(2x) was the correct answer but now I have three of them, what one is correct?

OpenStudy (anonymous):

But why didn't you find the derivative of y?

sam (.sam.):

No that was a mistake you should replace the "x" with "y"

OpenStudy (anonymous):

Even so, there is no 2y for the derivative, just y^2 meaning it wasn't derived.

OpenStudy (anonymous):

Anyone want to point out which one is correct? I'm seeing the 2/x-tanx as an answer elsewhere too